Grade 11Mechanics

An object moving with velocity u under acceleration a=∝t². find position of object after 2m/sec.

Riya sinha
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
 
Displacement is the double integeraation of accelration for continuos motion.
Therfore from the given data, 
a=∝t²
we can write it as:
a = kt^2
or
dv/dt = kt^2
dv = kt^2dt
Integerating both sides we get,
v = kt^3/3...............(2)
Again,
dx/dt = kt^3/3
or
dx = (kt^3/3)dt
Integerating both sides we get,
x = kt^4/12
At t=2 sec,
x = 4k/3 m.
This would be the displacement at the given time.
.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)